# Utility-Diversity Aware Online Batch Selection for LLM Supervised Fine-tuning

### v2

@@ 3.1 Intra-sample Importance Score via Nuclear Norm
Larger Nuclear Norm Positively Correlates With Higher Intra-sample Diversity.
Beyond optimization utility, a larger nuclear norm can also arise from the structural diversity of token-level output distributions within a sequence. Recall Lemma 3.1: for a fixed Frobenius norm, the nuclear norm achieves its minimum when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is rank-$1$ with only one singular value, and its maximum when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank with equal singular values. These two extremes correspond to distinct regimes of intra-sample diversity. The lower bound typically occurs when the model predicts only a single vocabulary token throughout the sequence (indicating repetitive generation behavior), while the upper bound is achieved when the model produces a florid and diverse vocabulary distribution across tokens (reflecting richer semantic information content). Figure 4 presents a simplified example illustrating what intra-sample diversity means in real cases. Intuitively, A larger nuclear norm implies that more vocabularies in this sequence are predicted and utilized, indicating higher prediction dispersity. We additionally provide an empirically strong correlation between $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$ and $rank(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t}))$ in Figure 3 to support this. Specific explanations are listed below.
Low Nuclear Norm Positively Correlates with Low rank and collinear rows (low diversity). Considering the lower bound of Lemma 3.1, since $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is rank-1, the row vectors $\bm{l}_{n}$ are nearly collinear, i.e., $\bm{l}_{n}\approx\alpha_{n}\bm{v}$ for some fixed $\bm{v}$. This degenerate spectral structure indicates minimal diversity: the model’s predictions for different tokens are aligned along the same direction, resulting in repetitive output.
High Nuclear Norm Positively Correlates with High rank and orthogonal rows (high diversity). Considering the higher bound of Lemma 3.1, $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank and all singular values are approximately equal. This occurs when the row vectors are diverse and orthogonal with comparable norms. This flat spectrum indicates maximal diversity: each token is semantically meaningful to the output, and predictions are spread across diverse directions, capturing a wide range of semantic information.
@@ 3.2 Inter-Sample Importance Score via Low-rank Similarity Matching
Current online batch selection methods (wang2024greats) only consider sample diversity within the candidate batch, which is suboptimal because the capacity of the batch is much smaller than the global datasets. To enhance global diversity, we maintain a fixed-size First-In-First-Out (FIFO) memory buffer $\bm{Q}\in\mathbb{R}^{M\times d}$, which stores representations $\bm{z}_{t}^{i}\in\mathbb{R}^{d}$ of the last $M$ samples selected for training ($M\gg B$), and measure inter-sample diversity by calculating the distance between the candidate sample and the recent training history.
@@
$s_{\text{inter}}^{t,i}=\frac{1}{|\bm{Q}|}\sum_{\bm{z}_{j}\in\bm{Q}}\|\bm{z}_{t}^{i}-\bm{z}_{j}\|_{2}.$ (9)
@@
As mentioned, the logits matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$ is semantically meaningful for computing diversity, but directly storing the whole matrix in the memory buffer is prohibitive (e.g., $\sim 74$GB for $1024$ samples in Qwen-2.5-7B). A natural strategy is to randomly project it into low-dimensional vectors $\bm{z}_{t}^{i}\in\mathbb{R}^{d}$, while preserving pairwise distances (johnson1984extensions). However, a direct projection using $\bm{\Gamma}\in\mathbb{R}^{NV\times d}$ also incurs severe storage cost for the down-projection matrix ($\sim 74$GB in Qwen-2.5-7B if $d$ is set to $1024$). To avoid this, we try to factorize $\bm{\Gamma}$ into two smaller projections: $\bm{\Gamma}_{1}\in\mathbb{R}^{d_{1}\times V}$ reducing the vocabulary dimension and $\bm{\Gamma}_{2}\in\mathbb{R}^{d_{2}\times N}$ reducing the sequence length dimension, which together approximate the original projection with $d=d_{1}d_{2}$. Formally, we aim to find proper $\bm{\Gamma}_{1}$ and $\bm{\Gamma}_{2}$ that operate in the following way:
@@
Fortunately, the subsampled randomized Fourier transform (SRFT)-style construction for $\bm{\Gamma}_{1}$ and $\bm{\Gamma}_{2}$ (see Theorem 3.2) can fulfill this role (ailon2006approximate; jin2021faster). This approach achieves an approximate Johnson-Lindenstrauss embedding while avoiding the storage of an explicit $NV\times d$ projection matrix and reduces the computational complexity from $\mathcal{O}(NVd)$ to $\mathcal{O}\left((N+V)d\log(NV)\right)$. The complete proof is provided in Appendix C.3.
@@
Let $N,V\in\mathbb{N}$, and choose $d_{1}\leq V$, $d_{2}\leq N$ with $d=d_{1}d_{2}$. Define
$\bm{\Gamma}_{1}=\sqrt{\frac{V}{d_{1}}}\bm{S}_{1}\bm{F}_{1}\bm{D}_{1}\in\mathbb{R}^{d_{1}\times V},\qquad\bm{\Gamma}_{2}=\sqrt{\frac{N}{d_{2}}}\bm{S}_{2}\bm{F}_{2}\bm{D}_{2}\in\mathbb{R}^{d_{2}\times N},$
where $\bm{F}_{1},\bm{F}_{2}$ are orthonormal transforms (discrete Fourier transform (DFT) matrices), $\bm{D}_{1},\bm{D}_{2}$ are random $\{\pm 1\}$ diagonal matrices with independent Rademacher entries on the diagonal, and $\bm{S}_{1},\bm{S}_{2}$ are random selection matrices that choose $d_{1}$ and $d_{2}$ rows uniformly at random without replacement. For each $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$, define
$\bm{u}_{t}^{i}=\operatorname{vec}(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t}))\in\mathbb{R}^{NV},\qquad\bm{v}_{t}^{i}=\operatorname{vec}(\bm{\Gamma}_{2}\cdot\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\cdot\bm{\Gamma}_{1}^{\top})\in\mathbb{R}^{d}.$
@@ 4 Experiments
4.1 Experimental Setup
Datasets and Backbones:
We evaluate online batch selection methods’ performance across four key domains: (1) general knowledge understanding using the MMLU (hendryckstest2021) benchmark, with auxiliary training datasets for fine-tuning and the official test set for evaluation; (2) scientific question answering using ScienceQA (lu2022learn) for both fine-tuning and evaluation; (3) mathematical reasoning on GSM8K (cobbe2021gsm8k) for both fine-tuning and evaluation; and (4) code generation using CodeAlpaca-20k (codealpaca) for training and HumanEval (chen2021evaluating) for evaluation. All experiments are conducted using Llama-3.1-8B (grattafiori2024llama) and Qwen-2.5-7B (yang2024qwen2) as backbone models following zou2025flylora.
Baselines:
We compare UDS against the following baselines: (1) Regular, where all samples are used without data selection; (2) Random Selection, which randomly chooses samples from batches; (3) MaxLoss (loshchilov2015online), which prioritizes samples with the highest training loss; (4) MaxGrad (katharopoulos2018not), which selects samples with the largest gradient norms; (5) RHO-Loss (mindermann2022prioritized), a representative method using a reference model; and (6) GREATS (wang2024greats), a state-of-the-art (SOTA) online batch selection approach. All baselines select the same fraction of data for a fair comparison.
Implementation Details:
Under hardware constraints, we employ LoRA (rank=8) for SFT training. The batch size is set to $B=8$ for all datasets. The optimal data selection ratio $\alpha$ is highly dependent on both the backbone model and the dataset; the ratios we adopt for different combinations of backbones and datasets are reported in Table 5. The default choice of the buffer size $M=1024$, the down-projection dimension is $d_{1}=128$, $d_{2}=8$, and we use this across all the experiments. For detailed sensitivity analysis, please refer to Appendix B.1. We report average accuracy on MMLU, ScienceQA, and GSM8K, and Pass@1 for HumanEval. We also compare training speed by throughput relative to an NVIDIA GeForce RTX 3090 GPU, following wang2024greats. All benchmarks use zero-shot evaluation and are repeated four times with different random seeds. More detailed experimental settings are provided in Appendix D.
4.2 Highest Accuracy with Lower Training Time than Full Dataset
@@
UDS achieves the highest accuracy.
UDS is more efficient than training on the full dataset.
4.3 Ablation Studies
@@
Both components of UDS contribute to performance.
We conduct ablation studies on the two components of UDS, as summarized in Table 4. Using only the Nuclear Norm consistently outperforms random selection, underscoring the importance of capturing intra-sample utility and diversity. Using only the Diversity Distance also improves performance over the baseline, as discouraging redundancy among selected samples enhances overall utility. Combined, the two components complement each other to achieve the best results across all benchmarks, highlighting the necessity of jointly modeling sample utility, intra-sample diversity, and inter-sample diversity.
@@ 4.4 Comparison across Different Data Scales
To assess the effectiveness of UDS under varying batch selection fractions, we conduct experiments as shown in Figure 5, where the horizontal axis denotes the number of selected samples $K$ per batch and the vertical axis reports average accuracy.
While most methods improve with $K$, UDS’s performance peaks at $K=4$, achieving the highest accuracy before gradually declines as additional, less informative samples are added. When $K=B=8$, all methods reduce to full-dataset fine-tuning. UDS consistently outperforms baselines, showing its ability to prioritize informative and diverse samples. At its peak, UDS even surpasses full-dataset fine-tuning, demonstrating that a small, well-curated subset can deliver both efficiency and accuracy. These results confirm that UDS effectively balances exploitation and exploration, yielding robust improvements under different selection budgets.
@@ Appendix A Related Work
Several studies have investigated online batch selection prior to the era of LLMs. Works like (jiang2019accelerating; katharopoulos2018not; loshchilov2015online) propose heuristic methods that select “important” samples (e.g., via high training loss, or large gradient norm). These methods are straightforward and often effective for smaller models or simpler tasks, but more recent studies (wang2024greats) show that these heuristics may not scale well for LLMs. Another paradigm uses a reference model to help evaluate sample importance. For instance, (deng2023towards; mindermann2022prioritized) use a pre-trained or separately trained reference model, and select samples based on how “different” they are relative to that model. These methods tend to require extra computation (e.g. extra forward passes) and sometimes held-out data to train the reference model, which can make them costly or impractical in large-scale LLM training. (kaddour2023no) points out that these overheads can outweigh benefits in many settings. GREATS (wang2024greats) is a more recent SOTA method. It selects data batches online by considering loss reduction on a validation set, and uses a “ghost inner-product” approximation (via a Taylor-expansion‐based greedy algorithm) to accelerate selection. However, the need for a validation set may be unavailable or non-representative in practical deployment (wang2024greats).
@@
Most existing approaches to data selection are performed in an offline manner, where samples are filtered prior to training (albalak2024survey; xia2024less; zhou2023lima; chen2023alpagasus; xie2023data; wettig2024qurating). This is primarily due to efficiency concerns, since curating the entire dataset at every iteration is prohibitively expensive. Early approaches extend uncertainty estimation and active learning heuristics, such as filtering by entropy or difficulty, but these remain limited in capturing truly informative samples (xu2020curriculum). A more impactful direction is large-scale pruning and deduplication: (lee2021deduplicating; tirumala2023d4) show that removing near-duplicate or low-quality data improves efficiency and prevents overfitting. In instruction tuning, several works highlight that “quality outweighs quantity”, with methods like LIMA (zhou2023lima), LESS (xia2024less), and Self-Instruct (wang2022self) filtering diverse, representative subsets while discarding noisy prompts. Recent advances further explore offline data selection through various sampling strategies: sachdeva2024train introduce LLM-as-judge method and density-based sampling; axiotis2024data propose clustering-based sensitivity sampling; and deb2025fishersft leverage information gain estimation to identify high-value examples. While effective to some extent, this perspective assumes that the value of each example is fixed throughout the learning process. In reality, learning is highly dynamic. Data that appear highly informative at the beginning of training may later become redundant (wang2024greats). This limitation motivates serving online batch selection as a complement for dynamically assessing sample value during the training process.
Appendix B Additional Results
B.1 Parameter Analysis

### v3

@@ 3.1 Intra-sample Importance Score via Nuclear Norm
Larger Nuclear Norm Correlates With Higher Intra-sample Diversity.
Beyond optimization utility, a larger nuclear norm can also arise from the structural diversity of token-level output distributions within a sequence. Recall Lemma 3.1: for a fixed Frobenius norm, the nuclear norm achieves its minimum when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is rank-$1$ with only one singular value, and its maximum when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank with equal singular values. These two extremes correspond to distinct regimes of intra-sample diversity. The lower bound typically occurs when the model predicts only a single vocabulary token throughout the sequence (indicating repetitive generation behavior), while the upper bound is achieved when the model produces a florid and diverse vocabulary distribution across tokens (reflecting richer semantic information content). Figure 3 presents a simplified example illustrating what intra-sample diversity means in real cases. Intuitively, A larger nuclear norm implies that more vocabularies in this sequence are predicted and utilized, indicating higher prediction dispersity. Figure 2(b) further supports this through the strong empirical correlation between $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$ and $\mathrm{rank}(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t}))$.
(i) Low Nuclear Norm Correlates with Low rank and collinear rows (low diversity). Considering the lower bound of Lemma 3.1, rank-1 $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ implies near collinear row vectors $\bm{l}_{n}\approx\alpha_{n}\bm{v}$ for some fixed $\bm{v}$. This degenerate spectral structure indicates minimal diversity: the model’s predictions for different tokens are aligned along the same direction, resulting in repetitive output.
(ii) High Nuclear Norm Correlates with High rank and orthogonal rows (high diversity). Considering the upper bound of Lemma 3.1, $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank and equal singular values. This occurs when the rows are diverse and orthogonal with comparable norms. This flat spectrum indicates maximal diversity: each token is semantically meaningful, and predictions are spread across diverse directions, capturing a wide range of semantic information.
@@ 3.2 Inter-Sample Importance Score via Low-rank Similarity Matching
Current online batch selection methods (Wang et al., 2024) only consider diversity within the candidate batch, which is suboptimal because the capacity of the batch is much smaller than the global datasets. To enhance global diversity, we maintain a fixed-size First-In-First-Out (FIFO) memory buffer $\bm{Q}\in\mathbb{R}^{M\times d}$, which stores representations $\bm{z}_{t}^{i}\in\mathbb{R}^{d}$ of the last $M$ samples selected for training ($M\gg B$), and measure inter-sample diversity by calculating the distance between the candidate sample and the recent training history.
@@
$s_{\text{inter}}^{t,i}=\underbrace{\frac{1}{|\bm{Q}|}\sum_{\bm{z}_{j}\in\bm{Q}}\|\bm{z}_{t}^{i}-\bm{z}_{j}\|_{2}}_{\text{diversity w.r.t. buffer}}+\underbrace{\frac{1}{|\mathcal{B}_{t}|}\sum_{\bm{z}_{k}\in\mathcal{B}_{t}}\|\bm{z}_{t}^{i}-\bm{z}_{k}\|_{2}}_{\text{(optional) intra-batch diversity}},$ (9)
@@
Remark: While within-batch diversity could theoretically improve performance, it is likely negligible since the batch size $B\ll M$ and data are shuffled. Similar samples rarely appear in a batch without being in the buffer. We thus omit this for simplicity. We provide an additional experiment in Appendix C.7 to validate this assumption.
@@
As mentioned before, the logits matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$ is semantically meaningful for computing diversity, but directly storing the whole matrix in the memory buffer is prohibitive (e.g., $\sim 74$GB for $1024$ samples in Qwen-2.5-7B). A natural strategy is to randomly project it into low-dimensional vectors $\bm{z}_{t}^{i}\in\mathbb{R}^{d}$, while preserving pairwise distances (Johnson et al., 1984). However, a direct projection using $\bm{\Gamma}\in\mathbb{R}^{NV\times d}$ also incurs severe storage cost for the down-projection matrix ($\sim 74$GB in Qwen-2.5-7B if $d$ is set to $1024$). To avoid this, we try to factorize $\bm{\Gamma}$ into two smaller projections: $\bm{\Gamma}_{1}\in\mathbb{R}^{d_{1}\times V}$ reducing the vocabulary dimension and $\bm{\Gamma}_{2}\in\mathbb{R}^{d_{2}\times N}$ reducing the sequence length dimension, which together approximate the original projection with $d=d_{1}d_{2}$. Formally, we aim to find proper $\bm{\Gamma}_{1}$ and $\bm{\Gamma}_{2}$ that operate in the following way:
@@
Fortunately, the subsampled randomized Fourier transform (SRFT)-style construction for $\bm{\Gamma}_{1}$ and $\bm{\Gamma}_{2}$ (see Theorem 3.2) can fulfill this role (Ailon and Chazelle, 2006; Jin et al., 2021). This approach achieves an approximate Johnson-Lindenstrauss embedding while avoiding the storage of an explicit $NV\times d$ projection matrix and reduces the computational complexity from $\mathcal{O}(NVd)$ to $\mathcal{O}\left((N+V)d\log(NV)\right)$. The complete proof is provided in Appendix D.3.
@@
Let $N,V\in\mathbb{N}$, and choose $d_{1}\leq V$, $d_{2}\leq N$ with $d=d_{1}d_{2}$. Define $\bm{\Gamma}_{1}\in\mathbb{R}^{d_{1}\times V}$ and $\bm{\Gamma}_{2}\in\mathbb{R}^{d_{2}\times N}$ as
$\bm{\Gamma}_{1}=\sqrt{\frac{V}{d_{1}}}\bm{S}_{1}\bm{F}_{1}\bm{D}_{1},\quad\bm{\Gamma}_{2}=\sqrt{\frac{N}{d_{2}}}\bm{S}_{2}\bm{F}_{2}\bm{D}_{2},$
where $\bm{F}_{1},\bm{F}_{2}$ are orthonormal transforms (discrete Fourier transform (DFT) matrices), $\bm{D}_{1},\bm{D}_{2}$ are random $\{\pm 1\}$ diagonal matrices with independent Rademacher entries on the diagonal, and $\bm{S}_{1},\bm{S}_{2}$ are random selection matrices that choose $d_{1}$ and $d_{2}$ rows uniformly at random without replacement. For each $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$, define $\bm{u}_{t}^{i}\in\mathbb{R}^{NV}$ and $\bm{v}_{t}^{i}\in\mathbb{R}^{d}$
$\bm{u}_{t}^{i}=\operatorname{vec}(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})),\quad\bm{v}_{t}^{i}=\operatorname{vec}(\bm{\Gamma}_{2}\cdot\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\cdot\bm{\Gamma}_{1}^{\top}).$
@@ 4 Experiments
4.1 Experimental Setup
Datasets and Backbones:
We mainly evaluate online batch selection methods’ performance across four key domains: (1) general knowledge understanding using the MMLU (Hendrycks et al., 2021a) benchmark, with auxiliary training datasets for fine-tuning and the official test set for evaluation; (2) scientific question answering using ScienceQA (Lu et al., 2022) for both fine-tuning and evaluation; (3) mathematical reasoning on GSM8K (Cobbe et al., 2021) for both fine-tuning and evaluation; and (4) code generation using CodeAlpaca-20k (Chaudhary, 2023) for training and HumanEval (Chen et al., 2021) for evaluation. All experiments are conducted using Llama-3.1-8B (Grattafiori et al., 2024) and Qwen-2.5-7B (Yang et al., 2024) as backbone models following (Zou et al., 2025).
Baselines:
We mainly compare UDS against the following online batch selection baselines: (1) Regular, where all samples are used without data selection; (2) Random Selection, which randomly chooses samples from batches; (3) MaxLoss (Loshchilov and Hutter, 2015), which prioritizes samples with the highest training loss; (4) MaxGrad (Katharopoulos and Fleuret, 2018), which selects samples with the largest gradient norms; (5) RHO-Loss (Mindermann et al., 2022), a representative method using a reference model; and (6) GREATS (Wang et al., 2024), a state-of-the-art (SOTA) online batch selection approach. All baselines select the same fraction of data for a fair comparison.
Implementation Details:
Under hardware constraints, we employ LoRA (rank=8) for SFT training. The batch size is set to $B=8$ for all datasets. The optimal data selection ratio $\alpha$ is highly dependent on both the backbone model and the dataset; the ratios we adopt for different combinations of backbones and datasets are reported in Table 5. The default choice of the buffer size $M=1024$, the down-projection dimension is $d_{1}=128$, $d_{2}=8$, and we use this across all the experiments. We report average accuracy on MMLU, ScienceQA, and GSM8K, and Pass@1 for HumanEval. We also compare training speed by throughput relative to an NVIDIA GeForce RTX 3090 GPU, following Wang et al. (2024). All benchmarks use zero-shot evaluation and are repeated four times with different random seeds. More detailed experimental settings are provided in Appendix E.
Figure 4: Impact of $d_{1}$, $d_{2}$, and $M$ on Qwen-2.5 performance and memory usage on MMLU. Green curves show average accuracy; blue curves show peak memory. Performance improves with increasing $d_{1}$, $d_{2}$, and $M$, while memory usage increases only slightly.
4.2 Highest Accuracy with Lower Training Time than Full Dataset
UDS achieves the highest accuracy.
UDS is more efficient than training on the full dataset.
Appendices C.3C.5 provide further evaluations on full SFT, larger batch sizes, instruction-tuning models, Long-CoT, and OOD datasets, validating the robustness of UDS.
4.3 Ablation Studies
Both components of UDS contribute to performance.
Table 4 summarizes the ablations on the two components of UDS. Nuclear Norm alone consistently outperforms random selection, underscoring the importance of capturing intra-sample utility and diversity. Similarly, Diversity Distance improves upon the baseline by reducing redundancy. Combined, the two components complement each other to achieve the best results across all benchmarks, highlighting the necessity of jointly modeling sample utility, intra-sample diversity, and inter-sample diversity.
Appendix C.2 provides additional ablations on buffer policies, distance aggregation, and random matrix constructions.
4.4 Hyperparameter Analysis
Figure 4 shows the key design of the bilinear down-projection and memory buffer incur no severe memory burden. Appendix C.1 further analyzes the impact of $M$, $d_{1}$, $d_{2}$ regarding accuracy and resources, and $\alpha$ regarding accuracy.
@@ 4.5 Comparison across Different Data Scales
To assess the effectiveness of UDS under varying batch selection fractions, we conduct experiments as shown in Figure 5. We report the number of selected samples $K$ per batch versus average accuracy across baselines.
While most methods improve with $K$, UDS’s performance peaks at $K=4$, achieving the highest accuracy before gradually declines as additional, less valuable samples are added. When $K=B=8$, all methods reduce to full-dataset fine-tuning. UDS consistently outperforms baselines, showing its ability to prioritize valuable and diverse samples. At its peak, UDS even surpasses full-dataset fine-tuning, demonstrating that a small, well-curated subset can deliver both efficiency and accuracy. These results confirm that UDS effectively balances exploitation and exploration, yielding robust improvements under different selection budgets.
@@ 5 Conclusion
Impact Statement
This work presents an algorithmic advancement intended to streamline the SFT process through efficient online batch selection. By enabling models to learn from high-utility, diverse samples without the overhead of full-dataset training or external reference models, we provide a pathway to optimize resource utilization in machine learning workflows. As a general-purpose optimization technique, the proposed framework operates within the boundaries of existing datasets and architectures; thus, it does not generate novel societal risks. The broader impact of the technology will be determined by the specific downstream tasks and data distributions applied by practitioners.
@@ Appendix A Related Work
Several studies have investigated online batch selection prior to the era of LLMs. Works like Jiang et al. (2019); Katharopoulos and Fleuret (2018); Loshchilov and Hutter (2015) propose heuristic methods that select “important” samples (e.g., via high training loss, or large gradient norm). These methods are straightforward and often effective for smaller models or simpler tasks, but more recent studies (Wang et al., 2024) show that these heuristics may not scale well for LLMs. Another paradigm uses a reference model to help evaluate sample importance. For instance, Deng et al. (2023); Mindermann et al. (2022) use a pre-trained or separately trained reference model, and select samples based on how “different” they are relative to that model. These methods tend to require extra computation (e.g. extra forward passes) and sometimes held-out data to train the reference model, which can make them costly or impractical in large-scale LLM training. (Kaddour et al., 2023) points out that these overheads can outweigh benefits in many settings. GREATS (Wang et al., 2024) is a more recent SOTA method. It selects data batches online by considering loss reduction on a validation set, and uses a “ghost inner-product” approximation (via a Taylor-expansion‐based greedy algorithm) to accelerate selection. However, the need for a validation set may be unavailable or non-representative in practical deployment (Wang et al., 2024).
@@
Most existing approaches to data selection are performed in an offline manner, where samples are filtered prior to training (Albalak et al., 2024; Xia et al., 2024; Zhou et al., 2023; Chen et al., 2023; Xie et al., 2023; Wettig et al., 2024). This is primarily due to efficiency concerns, since curating the entire dataset at every iteration is prohibitively expensive. Early approaches extend uncertainty estimation and active learning heuristics, such as filtering by entropy or difficulty, but these remain limited in capturing truly informative samples (Xu et al., 2020). A more impactful direction is large-scale pruning and deduplication: (Lee et al., 2021; Tirumala et al., 2023) show that removing near-duplicate or low-quality data improves efficiency and prevents overfitting. In instruction tuning, several works highlight that “quality outweighs quantity”, with methods like LIMA (Zhou et al., 2023), LESS (Xia et al., 2024), and Self-Instruct (Wang et al., 2022) filtering diverse, representative subsets while discarding noisy prompts. Recent advances further explore offline data selection through various sampling strategies: Sachdeva et al. (2024) introduce LLM-as-judge method and density-based sampling; Axiotis et al. (2024) propose clustering-based sensitivity sampling; and Deb et al. (2025) leverage information gain estimation to identify high-value examples. While effective to some extent, this perspective assumes that the value of each example is fixed throughout the learning process. In reality, learning is highly dynamic. Data that appear highly informative at the beginning of training may later become redundant (Wang et al., 2024). This limitation motivates serving online batch selection as a complement for dynamically assessing sample value during the training process.
Appendix B Discussions
B.1 Discussion on the Different Objectives of Offline Data Selection and Online Data Selection
Many data selection methods exist for SFT. However, most of them are conducted offline; that is, data selection is performed before model training begins. In this work, we focus on online batch selection for SFT. Offline and online selection methods serve different purposes and are not directly comparable, particularly in terms of training-time efficiency. Instead, they form a complementary pipeline: offline selection provides a coarse filtering stage to ensure dataset quality, while online selection adaptively captures the changing importance of samples throughout the training process, thereby improving overall efficiency and accuracy (Wang et al., 2024). In this sequential pipeline, the candidate data points for online batch selection are usually already carefully curated after offline data selection, from which “outliers, noisy data, or adversarial examples” have already been filtered. Nevertheless, even with a high-quality SFT dataset, the importance of samples can shift throughout the learning process (Wang et al., 2024). Compared to offline data selection, the core objective of online batch selection is to capture these dynamic changes in data importance during training, which is precisely what our UDS design addresses.
B.2 Discussion on the Role of Online Batch Selection from the Perspective of Curriculum Learning
The online batch selection setting can be viewed as a form of implicit curriculum learning, and many prior works on curriculum learning have shown the benefit of scheduling data in an “easy-to-hard” order (Wang et al., 2021). In this work, we define data utility as loss reduction. People may have the question of whether this criterion may prefer “easy” points for the model to learn quickly. Actually, if a data point can be learned sufficiently, that is indeed what we expect to select at this stage. For other samples that are more complex, even if they are more informative, if the model is not yet strong enough to effectively capture their patterns (i.e., they are not learnable), they may not be useful for learning at the current stage. Within the implicit curriculum learning framework, as the model becomes stronger, it gains the ability to capture more complex data points; at that point, these more difficult samples can then be selected for training.
B.3 Discussion on the Difference between Online Batch Selection and Active Learning
While online batch selection and active learning both target data efficiency, they differ in their core assumptions and objectives. Classical active learning typically relies on a relatively static model snapshot to select samples, and thus struggles to capture the highly dynamic nature of training: data points that are informative early on may quickly become redundant as the model evolves. In contrast, online batch selection is embedded directly into the optimization loop and explicitly conditions selection on the current training state, allowing sample importance to change over time.
A second key distinction lies in the selection criteria. Active learning methods are predominantly driven by uncertainty estimation, which is well aligned with improving predictive performance (Peng et al., 2024) but only indirectly related to gradient-based optimization. Online batch selection, however, must be more tightly coupled with optimization utility, since its immediate effect is to modulate gradient updates rather than to alter the composition of the training dataset. As a result, online batch selection prioritizes signals that reflect potential loss reduction and training dynamics, making it fundamentally different from classical active learning in both motivation and mechanism.
Appendix C Additional Results
C.1 Parameter Analysis